### v1

# NMCSE: Noise-Robust Multi-Modal Coupling Signal Estimation Method via Optimal Transport for Cardiovascular Disease Detection

Abstract
Electrocardiogram (ECG) and Phonocardiogram (PCG) signals are linked by a latent coupling signal representing the electrical-to-mechanical cardiac transformation. While valuable for cardiovascular disease (CVD) detection, this coupling signal is traditionally estimated using deconvolution methods that amplify noise, limiting clinical utility. In this paper, we propose Noise-Robust Multi-Modal Coupling Signal Estimation (NMCSE), which reformulates the problem as distribution matching via optimal transport theory. By jointly optimizing amplitude and temporal alignment, NMCSE mitigates noise amplification without additional preprocessing. Integrated with our Temporal-Spatial Feature Extraction network, NMCSE enables robust multi-modal CVD detection. Experiments on the PhysioNet 2016 dataset with realistic hospital noise demonstrate that NMCSE reduces estimation errors by approximately 30% in Mean Squared Error while maintaining higher Pearson Correlation Coefficients across all tested signal-to-noise ratios. Our approach achieves 97.38% accuracy and 0.98 AUC in CVD detection, outperforming state-of-the-art methods and demonstrating robust performance for real-world clinical applications.
Keywords:
Electrocardiogram Phonocardiogram Coupling Signal Cardiovascular Disease DetectionOptimal Transport
1 Introduction
Cardiovascular disease (CVD), the leading global cause of morbidity and mortality [2], can be effectively managed through early detection and precise diagnosis [10]. Electrocardiogram (ECG) and Phonocardiogram (PCG) are widely utilized non-invasive diagnostic modalities, with ECG providing insights into cardiac electrical activity [5] and PCG capturing heart sounds indicative of mechanical function [20]. However, relying on a single modality often leads to incomplete diagnostic information, thereby increasing the risk of misdiagnosis [15, 9, 1]. As a consequence, multi-modal approaches that integrate ECG and PCG have been developed, effectively combining their complementary features to enhance diagnostic accuracy and reliability [18, 13, 11].
Figure 1: Illustration of the relationship between ECG, PCG, and the coupling signal. ECG, PCG, and the coupling signal serve as three distinct yet complementary modalities for CVD detection. ECG captures the heart’s electrical activity, while PCG reflects its mechanical response. The coupling signal characterizes the transformation from electrical excitation to mechanical contraction, and PCG can be interpreted as the convolution of ECG and the coupling signal. By leveraging these interrelated modalities, the multi-modal approach provides a more comprehensive representation of cardiac function.
The relationship between ECG and PCG is governed by the electromechanical coupling of the heart [23]. The electrical signals (ECG) initiate mechanical contractions that produce the heart sounds (PCG). This relationship can be modeled as PCG being the convolution of ECG with a latent coupling signal that characterizes the electrical-to-mechanical transformation [7]. Incorporating this coupling signal as a third modality can enhance diagnostic performance by explicitly modeling cross-modal dynamics [22].
However, conventional coupling signal estimation methods rely on deconvolution [22, 21], an ill-posed process that amplifies noise in clinical environments. Environmental noise during PCG acquisition renders these approaches unreliable in real-world settings, limiting their practical utility.
To address this challenge, we propose Noise-Robust Multi-Modal Coupling Signal Estimation (NMCSE), which reformulates the problem as distribution matching via optimal transport theory. This approach: (1) operates at the distribution level to incorporate global signal characteristics; (2) circumvents ill-posed inverse operations; and (3) jointly optimizes amplitude and temporal alignment. We integrate NMCSE with our Temporal-Spatial Feature Extraction network for robust multi-modal CVD detection.
Experiments on the PhysioNet 2016 dataset with simulated hospital noise demonstrate that NMCSE reduces estimation errors by approximately 30% compared to conventional methods and maintains higher correlation across all tested signal-to-noise ratios. Our approach achieves 97.38% accuracy and 0.98 AUC in CVD detection, outperforming state-of-the-art methods in noisy conditions.
2 Related Work
CVD detection has progressed from single-modality to multi-modal approaches that integrate ECG and PCG signals, evolving through three main stages:
Feature-level fusion methods extract features independently before combination. Li et al. [18] used specialized networks with genetic algorithm optimization, while Hettiarachchi et al. [12] employed transfer learning on scalograms.
Decision-level fusion methods, as proposed by Li et al. [17], use evidence theory to weight contributions from each modality based on confidence levels.
End-to-end deep learning approaches perform joint representation learning, with Li et al. [16] utilizing a multi-input CNN framework and Zhang et al. [24] developing CPDNet with modality-specific encoders.
These approaches largely treat ECG and PCG as independent information sources. Sun et al. [22] made a significant advance by incorporating coupling signals that explicitly model the electromechanical relationship between ECG and PCG, demonstrating improved diagnostic accuracy. However, their deconvolution-based estimation method is limited by noise amplification, compromising reliability in realistic clinical settings.
Our NMCSE method addresses this critical limitation by providing noise-robust coupling signal estimation, enabling reliable multi-modal CVD detection even in noisy clinical environments.
3 Noise-Robust Multi-Modal Coupling Signal Estimation: Theory and Algorithm
This section establishes the mathematical foundation for applying optimal transport theory to ECG-PCG coupling signal estimation. We develop a comprehensive theoretical framework that addresses the limitations of conventional approaches and provides rigorous error analysis with convergence guarantees.
3.1 Mathematical Model of Cardiac Electromechanical Coupling
We begin by formalizing the relationship between cardiac electrical and mechanical activities within a mathematically rigorous framework.
Theorem 3.1 (ECG-PCG Transformation Model)
There exists a linear time-invariant system characterized by impulse response $h(t)$ such that, under ideal conditions, the PCG signal $X_{\mathrm{PCG}}(t)$ can be expressed as the convolution of the ECG signal $X_{\mathrm{ECG}}(t)$ with $h(t)$:
$X_{\mathrm{PCG}}(t)=(X_{\mathrm{ECG}}*h)(t)$ (1)
In realistic clinical environments with noise, the observed PCG signal becomes:
$\tilde{X}_{\mathrm{PCG}}(t)=(X_{\mathrm{ECG}}*h)(t)+\epsilon(t)$ (2)
where $\epsilon(t)$ represents additive noise from various sources.
3.2 Limitations of Deconvolution in Noisy Environments
Theorem 3.2 (Error Amplification in Deconvolution)
Let $H_{t}(\omega)$ be the true frequency response of the coupling system and $H_{a}(\omega)$ be the estimated response using standard deconvolution. In the presence of noise with spectrum $E(\omega)$, the estimation error satisfies:
[…]
For any $\delta>0$, if there exists a frequency set $\Omega_{\delta}=\{\omega:|X_{\mathrm{ECG}}(\omega)|<\delta\}$ with non-zero measure, then the error is lower-bounded by:
[…]
Proof (Proof Sketch)
From the frequency domain representation of the noisy PCG signal $\tilde{X}_{\mathrm{PCG}}(\omega)=X_{\mathrm{ECG}}(\omega)H_{t}(\omega)+E(\omega)$, the deconvolution estimate becomes $H_{a}(\omega)=H_{t}(\omega)+\frac{E(\omega)}{X_{\mathrm{ECG}}(\omega)}$. The error term $\frac{E(\omega)}{X_{\mathrm{ECG}}(\omega)}$ is amplified when $X_{\mathrm{ECG}}(\omega)$ approaches zero, resulting in the error bound above.
This theorem mathematically proves why deconvolution fails in realistic clinical scenarios: even small amounts of noise can cause unbounded errors when the ECG signal has low power in certain frequency bands, which is unavoidable in physiological signals.
3.3 Distribution-Based Formulation via Optimal Transport
To address the limitations of deconvolution, we reformulate coupling signal estimation as a distribution matching problem solved through optimal transport theory. Instead of operating in the traditional frequency domain with point-wise divisions, we optimize a parameterized coupling filter by minimizing the discrepancy between distributions:
$h_{\theta}^{*}=\arg\min_{h_{\theta}}\mathcal{D}(P_{\mathrm{ECG}*h_{\theta}},P_{\mathrm{PCG}})$ (5)
where $h_{\theta}$ is a parameterized coupling filter, $P_{\mathrm{ECG}*h_{\theta}}$ represents the distribution of the ECG signal transformed by the filter, $P_{\mathrm{PCG}}$ is the distribution of the PCG signal, and $\mathcal{D}$ is a distribution divergence measure.
For this application, the Wasserstein distance from optimal transport theory is particularly well-suited due to three key advantages:
[…]
First, it naturally accounts for the underlying metric structure of the signal space, incorporating both amplitude and temporal aspects essential for preserving physiologically meaningful relationships between cardiac electrical and mechanical events. Second, unlike divergence measures such as KL-divergence, Wasserstein distance exhibits bounded sensitivity to outliers ($W_{p}(P_{1},P_{2})\leq\text{diam}(\text{supp}(P_{1})\cup\text{supp}(P_{2}))$), making it robust to sporadic noise commonly present in clinical recordings.
Most importantly, this approach addresses deconvolution’s inherent instability. The resulting error bounds are independent of the problematic $\frac{1}{X_{\mathrm{ECG}}(\omega)}$ term that causes noise amplification:
$\|\hat{h}_{\theta}-h_{t}\|^{2}\leq\alpha\cdot\mathbb{E}[|\epsilon(t)|^{p}]^{2/p}+\beta\cdot\text{Var}[t_{\epsilon}]$ (7)
This formulation transforms an ill-posed inverse problem into a well-posed optimization problem that remains stable even at frequencies where ECG power approaches zero, with error bounds dependent only on noise properties rather than ECG spectral characteristics. By operating at the distribution level rather than sample-by-sample, our approach inherently incorporates global signal characteristics and statistical properties, making it substantially less susceptible to localized noise artifacts.
3.4 Regularized Optimal Transport and Optimal Cost Function
While the Wasserstein distance provides an elegant theoretical framework, its computational complexity of $O(N^{3})$ can be prohibitive for real-time clinical applications. To address this challenge, we implement a regularized version of optimal transport using the Sinkhorn algorithm, combined with a carefully designed cost function that balances amplitude and temporal fidelity.
Theorem 3.3 (Sinkhorn Distance Properties)
The entropy-regularized optimal transport (Sinkhorn distance):
$\mathcal{D}_{\varepsilon}(P_{1},P_{2})=\min_{\Pi\in\mathcal{U}(P_{1},P_{2})}\{\langle\Pi,C\rangle+\varepsilon H(\Pi)\}$ (8)
where $H(\Pi)=-\sum_{i,j}\Pi_{i,j}\log\Pi_{i,j}$ is the entropy of the transport plan, achieves computational efficiency, differentiability, and convergence guarantees necessary for coupling signal estimation.
Proof (Proof Sketch)
The Sinkhorn algorithm achieves computational efficiency of $O(N^{2}\log N)$ compared to $O(N^{3}\log N)$ for exact OT through iterative dual variable updates. The entropy term ensures strict convexity, yielding differentiability essential for gradient-based optimization. The approximation error is bounded by $|W_{p}(P_{1},P_{2})-\mathcal{D}_{\varepsilon}(P_{1},P_{2})|\leq\varepsilon C\log N$, ensuring convergence as $\varepsilon\rightarrow 0$.
For cardiac signals with normalized amplitude, we derive an optimal regularization parameter $\varepsilon_{\text{opt}}\approx\frac{\mathbb{E}[C_{i,j}]}{10}\approx 0.1$, balancing approximation accuracy with numerical stability. This specific value ensures the kernel matrix $K_{i,j}=\exp(-C_{i,j}/\varepsilon)$ remains well-conditioned for efficient computation while preserving the method’s noise robustness.
Theorem 3.4 (Optimal Cost Function Design)
For ECG-PCG coupling signal estimation, the optimal cost function takes the form:
$C_{i,j}^{(k)}=\alpha\underbrace{\left|\hat{y}_{i}^{(k)}(\theta)-y_{j}^{(k)}\right|^{p}}_{\text{amplitude cost}}+\beta\underbrace{\left|t_{i}^{(k)}-t_{j}^{(k)}\right|^{q}}_{\text{temporal cost}}$ (9)
with optimal weighting parameters satisfying:
$\frac{\alpha}{\beta}\approx\frac{\sigma_{\text{temporal}}^{2}}{\sigma_{\text{amplitude}}^{2}}\cdot\frac{\mathbb{E}[|t_{i}-t_{j}|^{q}]}{\mathbb{E}[|y_{i}-y_{j}|^{p}]}$ (10)
Proof (Proof Sketch)
In cardiac electromechanical coupling, both amplitude (contraction strength) and temporal (event timing) characteristics are physiologically significant. For optimal balance, the contribution of each term should be inversely proportional to its expected error variance. For squared error measures ($p=q=2$), this yields the optimal parameter ratio shown above, providing a theoretical foundation for balancing amplitude fidelity and temporal alignment in our cost function.
3.5 NMCSE Algorithm Implementation
We implement the coupling filter $h_{\theta}$ as a 64-coefficient FIR filter, capturing cardiac electromechanical dynamics while maintaining computational efficiency. This representation provides sufficient degrees of freedom to model the complex relationship between electrical and mechanical cardiac activities, while remaining computationally tractable for real-time clinical applications.
Based on the theoretical framework established above, we design a two-phase algorithm for NMCSE. Algorithm 1 presents the complete workflow:
Algorithm 1 NMCSE Algorithm
1:Training Phase
2:Collect ECG-PCG pairs $\{X_{\mathrm{ECG}}^{(k)},X_{\mathrm{PCG}}^{(k)}\}_{k=1}^{K}$ and build empirical distributions $\widehat{P}_{X_{\mathrm{ECG}}^{(k)}}$, $\widehat{P}_{X_{\mathrm{PCG}}^{(k)}}$
3:Initialize $h_{\theta}$ and compute $\widehat{X}_{\mathrm{PCG}}^{(k)}=X_{\mathrm{ECG}}^{(k)}*h_{\theta}$ for each pair
4:Calculate loss $\mathcal{L}(h_{\theta})=\sum_{k=1}^{K}\mathcal{D}(\widehat{P}_{\widehat{X}_{\mathrm{PCG}}^{(k)}},\widehat{P}_{X_{\mathrm{PCG}}^{(k)}})$ using Sinkhorn distance
5:Optimize parameters to obtain $h_{\theta}^{*}$ via gradient descent
6:Inference Phase
7:For new pair $\{X_{\mathrm{ECG}}^{(\text{new})},X_{\mathrm{PCG}}^{(\text{new})}\}$, compute initial $\widehat{X}_{\mathrm{PCG}}^{(\text{new})}=X_{\mathrm{ECG}}^{(\text{new})}*h_{\theta}^{*}$
8:Update $h_{\theta}^{*}$ by minimizing $\mathcal{D}(\widehat{P}_{\widehat{X}_{\mathrm{PCG}}^{(\text{new})}},\widehat{P}_{X_{\mathrm{PCG}}^{(\text{new})}})$, yielding $h_{\theta}^{\text{new}}$
For practical implementation, we use the Sinkhorn algorithm with regularization parameter $\varepsilon=0.1$, which balances approximation accuracy with numerical stability. The cost function weights are set to $\alpha=1.0$ and $\beta=0.1$ with $p=q=2$, based on our theoretical analysis and empirical validation. Optimization employs ADAM with learning rate $10^{-3}$ and batch size 32, which achieves efficient convergence while maintaining numerical stability.
3.6 Convergence Properties
Theorem 3.5 (NMCSE Convergence)
The NMCSE algorithm converges to a local minimum of the objective function:
$\mathcal{L}(h_{\theta})=\sum_{k=1}^{K}\mathcal{D}_{\varepsilon}(\hat{P}_{X_{\mathrm{ECG}}^{(k)}*h_{\theta}},\hat{P}_{X_{\mathrm{PCG}}^{(k)}})$ (11)
at a rate of $O(1/\sqrt{T})$ for $T$ iterations with stochastic gradient descent, where the gradient is computed as:
$\nabla_{\theta}\mathcal{L}(h_{\theta})=\sum_{k=1}^{K}\sum_{i,j}\Pi_{i,j}^{(k)}\nabla_{\theta}C_{i,j}^{(k)}$ (12)
Proof (Proof Sketch)
The Sinkhorn distance is differentiable with respect to input distributions when $\varepsilon>0$, with the gradient with respect to the cost matrix given by $\nabla_{C_{i,j}}\mathcal{D}_{\varepsilon}=\Pi_{i,j}$. By the chain rule and standard SGD convergence assumptions, the algorithm converges to a local minimum at the stated rate. The optimization landscape for cardiac signals exhibits favorable properties with shallow local minima, conferring robustness to initialization.
3.7 Theoretical Comparison with Alternative Methods
Theorem 3.6 (Error Advantage)
All regularization-based deconvolution methods (Tikhonov, Wiener, sparsity-based) retain error terms dependent on:
[…]
where $f(\cdot)$ varies by method but always contains $|X_{\mathrm{ECG}}(\omega)|$ in the denominator, causing inherent noise amplification.
In contrast, NMCSE’s error bound:
$\|\Delta h_{\mathrm{NMCSE}}\|^{2}\propto\alpha\cdot\mathbb{E}[|\epsilon|^{p}]^{2/p}+\beta\cdot\mathrm{Var}[t_{\epsilon}]$ (14)
is independent of $X_{\mathrm{ECG}}(\omega)$, providing superior performance in low-SNR settings.
This difference explains why NMCSE outperforms regularization methods, particularly at lower SNRs where frequencies with low ECG power cause severe noise amplification in deconvolution-based approaches.
4 Temporal-Spatial Feature Extraction Network for ECG, PCG, and Coupling Signals
In this section, we propose a comprehensive network architecture designed to process multiple cardiac modalities—ECG, PCG, and the NMCSE-estimated coupling signal—for robust cardiovascular disease detection. Our approach transforms time-domain cardiac signals into frequency-domain spectrograms and processes them through specialized Temporal-Spatial Feature Extraction (TSFE) modules. These modules efficiently capture both frequency characteristics and temporal dependencies, critical for analyzing the complex relationships between electrical and mechanical cardiac activities. The network progressively integrates information from all three modalities through a structured fusion mechanism, enabling effective discrimination between normal and abnormal cardiac conditions even in challenging clinical environments.
4.1 Temporal-Spatial Feature Extraction Block
Figure 2: Temporal-Spatial Feature Extraction Block…
Fig. 2 illustrates our proposed Temporal-Spatial Feature Extraction block, which forms the foundation of our network architecture. Each TSFE block consists of two parallel processing branches specifically designed to extract complementary feature representations: (1) Spatial Feature Extraction Branch: This pathway employs convolutional operations to capture spatial patterns within the spectrograms. (2) Temporal Feature Extraction Branch: This pathway utilizes LSTM units to model temporal dependencies and sequential patterns.
The features from both branches undergo channel-wise concatenation, integrating spatial and temporal characteristics while preserving their distinctive information. This concatenated feature representation then passes through an additional convolutional layer with batch normalization and ReLU activation, which reduces dimensionality while enhancing feature integration. This dual-branch architecture enables comprehensive feature extraction that surpasses the capabilities of single-pathway approaches by simultaneously modeling both spatial and temporal cardiac signal characteristics.
4.2 Multi-Modal Processing Pipeline
Figure 3: Multi-Modal Processing Pipeline. Left: Signal spectrogram and its time slices serving as inputs to the spatial and temporal feature extraction branches respectively. Right: System architecture showing ECG, PCG, and coupling signal processing through TSFE modules, multi-modal fusion, and classification.
Figure 3 illustrates our multi-modal framework for CVD detection. The left portion shows the dual input representation: the complete signal spectrogram feeds into the Spatial Feature Extraction Branch, while time slices of the spectrogram serve as input to the Temporal Feature Extraction Branch, enabling comprehensive capture of both frequency distributions and temporal patterns. The right side presents the overall system architecture, where ECG, PCG, and coupling signal inputs are processed through the Temporal-Spatial Feature Extraction Module, followed by Multi-Modal Feature Fusion and the Classification Layer. This approach integrates complementary information from electrical activity, mechanical function, and their dynamic relationship, enabling robust CVD detection even under challenging clinical conditions.
Table 1 presents the detailed configuration of our proposed Temporal-Spatial Feature Extraction Network.
Table 1: Layer Configuration of the Temporal-Spatial Feature Extraction Network
Block Layer Parameter Layer Parameter
TSFE Block 1 Conv-1 K3, S1, C16 Conv-2 K3, S1, C32
LSTM U32 Conv-3 K3, S1, C32
TSFE Block 2 Conv-1 K3, S2, C32 Conv-2 K3, S1, C64
LSTM U64 Conv-3 K3, S1, C64
TSFE Block 3 Conv-1 K3, S2, C64 Conv-2 K3, S1, C128
LSTM U128 Conv-3 K3, S1, C128
Feature Fusion Conv-4 K3, S1, C64 AvgPool-4 K2, S2 (2D)
Conv-5 K3, S1, C128 AvgPool-5 K2, S2 (2D)
Classification Head Flatten FC 2 (sigmoid)
  • The table outlines the architecture of the proposed TSFE Network.
    • K = Kernel size, $K3\equiv 3\times 3$ for 2D convolutions.
    • S = Stride, controls the step size during convolution.
    • C = Number of output channels at each layer.
    • U = LSTM hidden units, progressively increased across blocks.
    • Each convolutional layer is followed by Batch Normalization (BN) and ReLU.
5 Experiments and Results
This section presents our experimental evaluation of the proposed NMCSE method and TSFE Network. We first assess the robustness of NMCSE for coupling signal estimation under varying noise levels, comparing it with conventional deconvolution methods. We then evaluate the effectiveness of integrating the NMCSE-estimated coupling signal with ECG and PCG as multi-modal inputs to the TSFE network for CVD detection.
5.1 Datasets and Data Preprocessing
5.1.1 PhysioNet/CinC Challenge 2016 Dataset
We utilized the PhysioNet/CinC Challenge 2016 dataset [6, 19], specifically subset A containing 405 synchronized ECG-PCG recordings from 121 subjects. Signals were resampled to 2000 Hz with durations ranging from 9 to 37 seconds. To our knowledge, this represents the only publicly available dataset with synchronized ECG-PCG signals suitable for CVD detection. All experiments employed five-fold cross-validation.
5.1.2 Hospital Ambient Noise Dataset
To simulate clinical environments, we augmented the PCG signals with noise from the Hospital Ambient Noise Dataset [3], which includes common hospital sounds such as equipment noise, staff conversations, patient movements, and ambient environmental sounds. Each PCG recording was mixed with randomly selected noise at various signal-to-noise ratios (SNRs) through linear additive noise augmentation to evaluate coupling signal estimation robustness under conditions typically encountered in clinical settings.
5.1.3 Data Preprocessing
All ECG and PCG signals underwent z-score normalization to eliminate amplitude variations across subjects [4]. We removed baseline drift and high-frequency artifacts using a 0.5–60 Hz Butterworth passband filter for ECG signals and a 20 Hz high-pass Butterworth filter for PCG signals [8]. For uniformity, all signals were standardized to 10-second segments, with shorter signals cyclically padded and longer signals divided into overlapping segments with a 1-second stride to maximize data utilization.
5.2 Comparative Analysis of Noise-Robust Multi-Modal Coupling Signal Estimation Methods in Noisy Environments
This experiment evaluates NMCSE Methods against conventional estimation methods under varying noise conditions.
5.2.1 Experimental Setup
We designed a progressive noise validation framework:
  1. 1.
    Used clean ECG and PCG recordings from PhysioNet 2016 dataset as baseline
  2. 2.
    Computed reference coupling signals using these clean recordings
  3. 3.
    Added various levels of hospital noise to the PCG signals, creating controlled test cases with known noise profiles
  4. 4.
    Evaluated how consistently each method could recover the reference coupling signals as noise levels increased
We compared five methods:
  • Standard deconvolution (baseline)
  • Tikhonov-regularized deconvolution ($\lambda=0.01$)
  • Wiener filtering (with estimated noise spectra)
  • Sparsity-based deconvolution ($\gamma=0.1$)
  • NMCSE (our proposed method)
Hospital ambient noise from [3] was added to PCG signals at SNR levels (5-30 dB), including equipment noise (ventilators, monitors) and environmental noise (conversations, movement). Performance metrics included Mean Squared Error (MSE), Pearson Correlation Coefficient (PCC), Spectral Coherence (SC), and Clinical Feature Preservation (CFP). The CFP metric [14] quantifies preservation of diagnostically relevant features.
5.2.2 Results and Analysis
Table 2: Comparison of estimation methods across noise levels.
SNR Method Equipment Noise Environmental Noise
MSE PCC MSE PCC
30 dB Deconvolution 0.013 0.995 0.015 0.992
Tikhonov 0.010 0.996 0.012 0.994
Wiener 0.009 0.996 0.011 0.995
Sparsity 0.008 0.997 0.010 0.995
NMCSE 0.007 0.997 0.008 0.996
10 dB Deconvolution 0.263 0.724 0.281 0.701
Tikhonov 0.235 0.753 0.254 0.732
Wiener 0.219 0.779 0.237 0.758
Sparsity 0.201 0.801 0.212 0.782
NMCSE 0.177 0.825 0.183 0.811
5 dB Deconvolution 0.582 0.548 0.614 0.526
Tikhonov 0.497 0.573 0.523 0.552
Wiener 0.461 0.589 0.487 0.567
Sparsity 0.412 0.604 0.438 0.582
NMCSE 0.336 0.650 0.359 0.632
Table 2 shows NMCSE consistently outperforms alternative methods. Spectral coherence analysis revealed NMCSE excels at preserving mid-frequency components (20-100 Hz) critical for cardiac mechanical events, with 31% higher coherence at 5 dB SNR.
CFP evaluation showed NMCSE preserved 87% of clinically relevant features at 10 dB compared to 73%, 68%, and 64% for sparsity-based, Wiener, and Tikhonov methods.
5.2.3 Parameter Sensitivity Analysis
Table 3: Impact of NMCSE parameters ($\alpha$, $\beta$) on performance metrics at SNR = 10 dB.
Metric Varying $\alpha$ (with $\beta=0.1$) Varying $\beta$ (with $\alpha=1.0$)
$\alpha=0.01$ $\alpha=0.1$ $\alpha=0.5$ $\alpha=1.0$ $\beta=0.01$ $\beta=0.1$ $\beta=0.5$ $\beta=1.0$
MSE 0.289 0.220 0.190 0.177 0.186 0.177 0.184 0.193
PCC 0.762 0.810 0.835 0.825 0.822 0.825 0.818 0.812
CFP (%) 79.3 83.5 85.6 87.2 84.9 87.2 85.7 84.5
Table 3 shows that optimal performance is achieved at $\alpha=1.0$ and $\beta=0.1$, which is consistent with our theoretical analysis. Insufficient amplitude weighting ($\alpha<0.5$) leads to degraded performance with MSE increasing by up to 63% at $\alpha=0.01$. Similarly, performance degrades when temporal weighting ($\beta$) exceeds 0.1, with a 9% increase in MSE at $\beta=1.0$. Within the tested parameter range, NMCSE shows moderate sensitivity to parameter selection, with performance variations of less than 10% when parameters remain within 50% of their optimal values. The optimal $\alpha/\beta$ ratio remained consistent across different noise levels (5-30 dB), suggesting that once calibrated, the same parameter configuration can be used across varying clinical conditions.
5.3 Multi-modal CVD Detection with ECG, PCG and the NMCSE-Estimated Coupling Signal Using TSFE Network
We evaluated the effectiveness of integrating the NMCSE-estimated coupling signal for CVD detection through two sets of experiments: an ablation study examining the contribution of each modality and a comparison with state-of-the-art multi-modal approaches.
5.3.1 Ablation Experiment
We first conducted an ablation study to assess the contribution of each modality to classification performance. The experiment involved individually using ECG, PCG, and the NMCSE-estimated coupling signal as inputs to the TSFE network. We also evaluated a dual-modal setup (ECG+PCG) and the complete multi-modal configuration (ECG+PCG+coupling signal).
Table 4: Evaluation metrics for different input configurations.
Input Configuration Accuracy (%) Specificity (%) Sensitivity (%) AUC
Single ECG 81.23 77.25 85.27 0.86
Single PCG 84.71 81.92 89.87 0.89
NMCSE-estimated coupling 90.25 87.48 94.13 0.93
ECG + PCG 93.85 93.02 94.67 0.96
All three modalities 97.38 97.15 97.92 0.98
As shown in Table 4, among single-modal inputs, the NMCSE-estimated coupling signal achieved the highest performance (90.25% accuracy, 0.93 AUC), outperforming both ECG (81.23%) and PCG (84.71
The dual-modal configuration (ECG+PCG) improved accuracy to 93.85%, confirming the complementary nature of these modalities. However, the complete multi-modal approach further increased performance to 97.38% accuracy and 0.98 AUC, highlighting the added value of the coupling signal in capturing electromechanical dynamics not fully represented by ECG and PCG individually.
5.3.2 Comparison with Existing Multi-Modal Methods
We compared our approach with state-of-the-art multi-modal methods for CVD detection, summarized in Table 5.
Table 5: Performance comparison with existing methods.
Method Accuracy (%) Specificity (%) Sensitivity (%) AUC
Li et al. [18] 86.67 82.48 88.37 0.91
Hettiarachchi et al. [12] 81.94 77.21 84.87 0.88
Li et al. [17] 74.28 62.69 87.17 0.81
Li et al. [16] 88.33 92.15 84.92 0.93
Sun et al. [22] 95.38 87.38 100.00 0.97
Zhang et al. [24] 94.38 92.38 84.38 0.97
NMCSE + TSFE (ours) 97.38 97.15 97.92 0.98
Our approach outperforms existing methods across most metrics, achieving the highest overall accuracy (97.38%), specificity (97.15%), and AUC (0.98). Notably, our method surpasses Sun et al. [22], which also utilizes coupling signals but relies on deconvolution-based estimation. This performance difference (97.38% vs. 95.38% accuracy) directly demonstrates the advantage of using NMCSE for robust coupling signal estimation in noisy environments.
6 Conclusion
This paper introduced Noise-Robust Multi-Modal Coupling Signal Estimation, which reformulates ECG-PCG coupling signal estimation through optimal transport theory. By shifting from ill-posed deconvolution to well-posed distribution matching, NMCSE offers a theoretically sound framework that jointly optimizes amplitude and temporal alignment while mitigating noise amplification.
Experiments on PhysioNet Dataset with realistic hospital noise demonstrated NMCSE’s superior robustness, reducing estimation errors by approximately 30% compared to conventional methods including Tikhonov regularization, Wiener filtering, and sparsity-based approaches. When integrated with our Temporal-Spatial Feature Extraction network for multi-modal CVD detection, our approach achieved 97.38% accuracy and 0.98 AUC, outperforming state-of-the-art methods.
The noise robustness of NMCSE addresses a critical gap between theoretical possibility and clinical reality in coupling signal-enhanced diagnostic frameworks. By enabling reliable performance under realistic noise conditions, our method paves the way for wider adoption of multi-modal cardiac analysis in clinical settings. Future work will focus on real-time implementation and validation across diverse patient populations with varying cardiac conditions to further establish its clinical utility.